Weather Report for February 20, 2012

The unsettled weather that has been a hallmark of this observing campaign continued tonight. Clouds and thick cirrus forced a stop to work only about an hour after opening the dome at the start of the night. Only one early evening target was successfully observed. The nearly 6-hour halt to observing ended when the sky cleared at about 0130 local time and several more blazars could be observed during the remaining time before dawn. Seeing varied between 1.2 and 2 arcsec and the wind finally died down after midnight.